OpenStudy (anonymous):

Find the normal force exerted on a 2.3kg book resting on a surface inclined at 34∘ above the horizontal.

OpenStudy (shamim):

Can u draw it?

OpenStudy (shamim):

Anyway, u know the weight of ur book is=mg=2.3*9.8=? Acting directly downward. Right?

OpenStudy (shamim):

Component of weight perpendicular to the plane=2.3*9.8*cos34=?

OpenStudy (shamim):

U know every action has an equal and oposit reaction.

OpenStudy (shamim):

So 2.3*9.8*cos34 is a force acting by book on the plane surface. And this is action

OpenStudy (shamim):

So reaction also will b =2.3*9.8*cos34N=. It will b a force on the book by surface. And this is called normal reaction. Right?

OpenStudy (shamim):

Response plz!!!!!!